OpenStudy (anonymous):

The test to detect the presence of a liver disorder is 98% accurate for a person who has the disease and 97% accurate for a person who does not have the disease. If 3.5% of the people in a given population actually have the disorder, what is the probability that a randomly chosen person tests positive? 0.0343 0.035 0.06325 0.02895

OpenStudy (anonymous):

it is a lot easier to do this with numbers than with percents, although this is a probability question and we can use probabilities to solve it

OpenStudy (anonymous):

lets say you have 1,000 people since \(3.5\%\) have the disease, that means 35 people have it and therefore the rest, \(965\) do not have it

OpenStudy (anonymous):

but then what about the 97% and 98%?

OpenStudy (anonymous):

of the 35 who have the disease, 98% will test positive, so \(.98\times 35=34.3\) test positive of the 965 who do not have it, 97% will test negative, and 3% will test positive 3% of 965 is \(.03\times 965=28.95\) test positive

OpenStudy (anonymous):

the total number that test positive is therefore \(34.3+28.95=63.25\)

OpenStudy (anonymous):

then divide by 1000 to get your probability

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw